OpenStudy (anonymous):

plz help wll fan and metal plz help me Use estimation to check the answer to this story problem. The zoo has 332 butterflies in the enclosure. Twelve of the butterflies are yellow and 106 of the butterflies are red. Half of the remaining butterflies are orange. How many butterflies are orange? A student answered that there were 107 orange butterflies. Which statement about the student's estimation is correct? A. The answer should be about 280. The student's answer is not correct. B. The answer should be about 220. The student'

OpenStudy (anonymous):

Use estimation to check the answer to this story problem. The zoo has 332 butterflies in the enclosure. Twelve of the butterflies are yellow and 106 of the butterflies are red. Half of the remaining butterflies are orange. How many butterflies are orange? A student answered that there were 107 orange butterflies. Which statement about the student's estimation is correct? A. The answer should be about 280. The student's answer is not correct. B. The answer should be about 220. The student's answer is not correct. C. The answer should be about 200. The student's answer is not correct. D. The answer should be about 110. The student's answer is close to the estimated answer.

OpenStudy (stormer15):

@moutaindew Have you tried solving?

OpenStudy (anonymous):

before yes its hard for me tho

OpenStudy (stormer15):

Take it step by step 332 - 12 (yellow) = ?

OpenStudy (anonymous):

320

OpenStudy (stormer15):

320 - 106 (red) = ?

OpenStudy (anonymous):

214?

OpenStudy (stormer15):

214 / 2 = orange

OpenStudy (anonymous):

212

OpenStudy (stormer15):

no

OpenStudy (anonymous):

107

OpenStudy (stormer15):

which answer choice goes with that?

OpenStudy (anonymous):

d

OpenStudy (stormer15):

so D is ur answer =]